### v2

# On $L^{p}$-semigroup to Stokes equation with dynamic boundary condition in the half-space

Abstract.
We consider evolutionary Stokes system, coupled with the so-called dynamic boundary condition, in the simple geometry of a $d$-dimensional half-space. Using the Fourier transform, we obtain an explicit formula for the resolvent. Maximal regularity estimates and existence of analytic semigroup in the $L^{p}$-setting are then deduced using the methods of $\mathcal{H}^{\infty}$-calculus.
Key words and phrases:
Stokes equations, dynamic boundary conditions, unbounded domains, maximal regularity, analytic semigroup
2000 Mathematics Subject Classification:
76D07, 47D03, 35D65
The authors thank the Czech Science Foundation, Grant Number 20-11027X, for its support.
* Corresponding author.
1. Introduction
This paper is motivated by the system
(1) $\partial_{t}\bm{u}-\operatorname{div}2\nu D\bm{u}+\nabla\pi$ $=\bm{f},\qquad\operatorname{div}\bm{u}=0\quad\textrm{ in $(0,T)\times\Omega$},$
$\beta\partial_{t}\bm{u}+\alpha\bm{u}+2\nu[{(D\bm{u})}\bm{n}]_{\tau}$ $=\beta\bm{h},\qquad\bm{u}\cdot\bm{n}=0\quad\textrm{ on $(0,T)\times\Gamma$},$
where $\Omega=\{x_{d}>0\}$ is the $d$-dimensional half-space, and $\Gamma=\{x_{d}=0\}$ the corresponding boundary. Note that (1) consists of two evolution equations, coupled via the trace operator. The problem can be efficiently treated in the Hilbert $L^{2}$-setting, since there is an underlying symmetric operator $A:V\to V^{*}$, where $V$ is a suitable subspace to $W^{1,2}(\Omega)$, see e.g. [1], cf. also the weak formulation (6) below. However, for $p\neq 2$, the $L^{p}$-theory is expectedly more difficult and up to our knowledge, is so far missing in the literature. In the present paper, we intend to partially amend this, with two main results being: the existence of an analytic semigroup in $L^{p}(\Omega)\times L^{p}(\Gamma)$, and the optimal $W^{2,p}$ and $W^{1,p}$-regularity estimates for the corresponding stationary (resolvent) problem.
Although our results are formulated for the case of half-space $\Omega=\{x_{d}>0\}$ only, extension to more general domains should be possible, using standard localization and perturbation arguments. Note that this would require extension of our theorems to the case of functions $\bm{u}$ with non-zero divergence, which is rather straighforward, using the estimates of [8].
The paper is organized as follows: In Section 2, we introduce function spaces and state our main results: Theorem 2.1 and Theorem 2.2, providing resolvent estimates and optimal regularity in the $L^{p}$-setting. In Corollary 2.1, we show the existence of an analytic semigroup. We also identify the domain of its generator. Main technical results of the paper then follow: in Section 3, we use (partial) Fourier transform, in the spirit of [8], to express the resolvent in terms of the boundary right-hand side. Related key estimates in $L^{p}$ spaces are then obtained in Section 4. Finally, proofs of the main theorems are completed in Section 5, essentially by combining results of previous section with the well-known $L^{p}$-theory for the (homogeneous) Dirichlet problem.
Let us conclude this introduction by discussing some related, both classical and recent results. For the Stokes system subject to a homogeneous Dirichlet boundary conditions, the existence of $L^{p}$-analytic semigroup was proven in [10] via the theory of pseudo-differential operators. For a more elementary approach based on the Fourier transform and the classical multiplier theorems, see [8]. Recently, a different method using the maximal operator was proposed in [16]; however, the method seems to be limited to bounded domains.
There is also an extensive literature devoted to problems related to (1), but crucially, without the time derivative in the boundary equation. For example, a fairly complete $L^{p}$-theory, including the existence of an analytic semigroup, has been established in [2], see also [9] and the references therein; yet the results again seem to be limited to bounded $\Omega$ only. In the case of a half-space, the problem was treated in [15], albeit for $\bm{h}=0$ only. A more modern approach based on the $\mathcal{H}^{\infty}$-calculus was employed here.
A rather comprehensive functional setup for treating resolvent problems related to dynamic boundary conditions is developed in [7], see also the references therein. This method does provide sharp regularity estimates in fractional Sobolev spaces, yet direct application to Stokes problem – due to the incompressibility condition and the pressure term – does not seem viable.
Let us also mention recent paper [6], where the existence of an analytic semigroup for (1) is proven in case of smooth bounded domains, in the Hilbert setting of the maximal regularity space $L^{2}(\Omega)\times W^{1/2,2}(\Gamma)$.
Concerning possible applications, we remark that optimal regularity estimates of the stationary Stokes problem (cf. Theorem 2.2) can be used to obtain higher regularity of evolutionary nonlinear problems, via the standard bootstrap argument. See [14] for the corresponding analysis in bounded domains, where the smoothness of solutions was used to estimate dimension of the global attractor, in the case of the 2d Navier-Stokes equations with dynamic boundary conditions.
2. Notation and main results
We set $\Omega=\{(x^{\prime},x_{d})\in\mathbb{R}^{d-1}\times\mathbb{R};x_{d}>0\}$ and identify $\Gamma=\partial\Omega$ with $\mathbb{R}^{d-1}$. Furthermore, $\bm{n}=(0,-1)$ is the outer normal, and $\tau:\mathbb{R}^{d}\to\mathbb{R}^{d-1}$ denotes the tangential component. For $p\in(1,\infty)$ and $k\geq 1$ integer, we denote by $L^{p}$ and $W^{k,p}$ the usual Lebesgue and Sobolev spaces. We also need Besov spaces $B^{s,p}_{p}$ with a general real $s$; we allow for non-integer $s$ in the spirit of [4]. The subscripts $\sigma$ or $n$ stand for the solenoidality, and zero normal component, respectively. Finally, we will extensively use the partial Fourier transform $\mathcal{F}_{x^{\prime}\to\xi}$, which is defined by
$\mathcal{F}_{x^{\prime}\to\xi}\bm{u}(x^{\prime},x_{d})=\widehat{\bm{u}}(\xi,x_{d})=\int_{\mathbb{R}^{d-1}}\bm{u}(x^{\prime},x_{d})e^{-ix^{\prime}\cdot\xi}\,{\rm d}x^{\prime},\quad\xi\in\mathbb{R}^{d-1}.$
Since it will not play any role in our paper, we set $\beta=\nu=1$ in (1). We will be concerned with the resolvent problem related to it, i.e.
(2) $(\lambda-\Delta)\bm{u}+\nabla\pi$ $=\bm{f}\quad\text{ in }\Omega,$
(3) $\text{div}\,\bm{u}$ $=0\quad\text{ in }\Omega,$
(4) $\bm{u}\cdot\bm{n}$ $=0\quad\text{ on }\Gamma,$
(5) $(\lambda+\alpha)\bm{u}+2[(D\bm{u})\bm{n}]_{\tau}$ $=\bm{h}\quad\text{ on }\Gamma,$
where $\lambda$ is a complex number. Note that $\operatorname{div}2D\bm{u}=\Delta\bm{u}$. Multiplying (2) by $\bm{\varphi}\in C^{\infty}_{0,\sigma}(\overline{\Omega})$ and integrating by parts, one obtains the integral formulation
(6) $\begin{gathered}\lambda\int_{\Omega}\bm{u}\cdot\bm{\varphi}\,{\rm d}x+(\lambda+\alpha)\int_{\Gamma}\bm{u}\cdot\bm{\varphi}\,{\rm d}S+\int_{\Omega}2D\bm{u}:D\bm{\varphi}\,{\rm d}x\\ =\int_{\Omega}\bm{f}\cdot\bm{\varphi}\,{\rm d}x+\int_{\Gamma}\bm{h}\cdot\bm{\varphi}\,{\rm d}S.\end{gathered}$
This problem is easy to handle in the $L^{2}$-setting, using Korn’s inequality, see [2] or [1]. However, our objective is to obtain the $L^{p}$ theory, as stated in the following two main theorems.
Theorem 2.1.
Let $\alpha\geq 0$, $p\in(1,+\infty)$, $\bm{f}\in L^{p}_{\sigma,n}(\Omega)$ and $\bm{h}\in L^{p}_{n}(\Gamma)$. Then for any complex $\lambda$ with $\mathfrak{Re}\,\lambda>0$ there exists unique $\bm{u}\in W^{1,p}_{\sigma,n}(\Omega)$ a weak solution to (2)–(5).
If moreover $|\lambda|>\omega$ for some $\omega>0$ fixed, then
(7) $\lVert\bm{u}\rVert_{L^{p}(\Omega)}+\lVert\bm{u}\rVert_{L^{p}(\Gamma)}\leq\frac{C_{0}}{|\lambda|}\big{(}\lVert\bm{f}\rVert_{L^{p}(\Omega)}+\lVert\bm{h}\rVert_{L^{p}(\Gamma)}\big{)},$
where the constant $C_{0}$ depends on $d$, $p$ and $\omega$, but is independent of $\lambda$ and $\alpha$.
Theorem 2.2.
Let $\alpha\geq 0$, $p\in(1,+\infty)$ and $\lambda$ is complex with $\mathfrak{Re}\,\lambda>0$.
  • (i)
    Let $\bm{f}\in L^{r}_{\sigma,n}(\Omega)$, where $r$ is given by the relation $1/r=1/p+1/d$, and let $\bm{h}\in B^{-1/p,p}_{p;n}(\Gamma)$. Then one has the estimate
    (8) $\lVert\bm{u}\rVert_{W^{1,p}(\Omega)}\leq C_{1}\big{(}\lVert\bm{f}\rVert_{L^{r}(\Omega)}+\lVert\bm{h}\rVert_{B^{-1/p,p}_{p}(\Gamma)}\big{)}.$
  • (ii)
    Let $\bm{f}\in L^{p}_{\sigma,n}(\Omega)$ and $\bm{h}\in B^{1-1/p,p}_{p;n}(\Gamma)$. Then the solution to resolvent problem (2)-(5) is strong, and one has the estimate
    (9) $\lVert\bm{u}\rVert_{W^{2,p}(\Omega)}+\lVert\nabla\pi\rVert_{L^{p}(\Omega)}$ $\leq C_{2}\big{(}\lVert\bm{f}\rVert_{L^{p}(\Omega)}+\lVert\bm{h}\rVert_{B^{1-1/p,p}_{p}(\Gamma)}\big{)}.$
The constants $C_{1}$, $C_{2}$ depend on $d$, $p$ and $|\lambda|$.
Postponing proofs of these theorems to Section 5, we can now establish the existence of analytic semigroup.
Corollary 2.1.
System (1) generates an analytic semigroup in $L^{p}(\Omega)\times L^{p}(\Gamma)$, where the domain of generator is contained in $B^{1+1/p-\varepsilon,p}_{p}(\Omega)$, $\varepsilon>0$ arbitrary.
Proof.
Set $X=L^{p}_{\sigma,n}(\Omega)\times L^{p}_{n}(\Gamma)$, and define $A\bm{u}=(\bm{f},\bm{h})$ by the (weak) formulation of the resolvent, i.e. (6). We set
$\mathcal{D}(A)=\{\bm{u}\in W^{1,p}_{\sigma,n}(\Omega);\ A\bm{u}\in X\}$
– with a slight abuse of notation, we identify the Sobolev function $\bm{u}$ with the couple $(\bm{u},\operatorname{tr}\bm{u})$. It is easy to verify that $A$ is densely defined, and the closedness follows from the resolvent estimate (7); see also (43) below. By a standard result (see e.g. [5, Corollary 3.7.14]), the existence of semigroup follows.
Let us identify the domain of $A$. In view of uniqueness (cf. the argument after (43) below), we can treat the regularity w.r. to $\bm{f}\in L^{p}(\Omega)$ and $\bm{h}\in L^{p}(\Gamma)$ separately. If $\bm{h}=0$, if follows from (9) that $\bm{u}\in W^{2,p}(\Omega)$. However, in contrast to the Dirichlet problem, this cannot be the domain since (5) would then imply $\bm{h}\in B^{1-1/p,p}_{p}(\Gamma)$.
In fact, we claim that for any $\varepsilon>0$
(10) $B^{1+1/p+\varepsilon,p}_{p;\sigma,n}(\Omega)\subset\mathcal{D}(A)\subset B^{1+1/p-\varepsilon,p}_{p;\sigma,n}(\Omega).$
Indeed, interpolating (98) with $\bm{f}=0$, we get the second inclusion. On the other hand, if $\bm{u}\in B^{1+1/p+\varepsilon,p}_{p}(\Omega)$, then the left-hand side of (5) belongs to $B^{\varepsilon,p}_{p}(\Gamma)$, in contradiction to the fact that in general, $\bm{h}\in L^{p}(\Gamma)$ only.
[…]
3. Fundamental solution
Here, we deal with (2)-(5) for $\bm{f}=0$, $g=0$, i.e. we consider the system
(11) $(\lambda-\Delta)\bm{u}+\nabla\pi$ $=0\quad\text{ in }\Omega,$
(12) $\text{div}\,\bm{u}$ $=0\quad\text{ in }\Omega,$
(13) $\bm{u}\cdot\bm{n}$ $=0\quad\text{ on }\Gamma,$
(14) $(\lambda+\alpha)\bm{u}+2[(D\bm{u})\bm{n}]_{\tau}$ $=\bm{\phi}\quad\text{ on }\Gamma.$
We will (formally) solve the problem, using Fourier analysis.
3.1. Reducing the equations
By $\text{div}^{\prime}$, $\Delta^{\prime}$ we denote the corresponding operator with respect to $x^{\prime}$. Let us now simplify (13)-(14). First, $\bm{n}=(0,-1)$ and thus
(15) $\bm{u}\cdot\bm{n}=0\Leftrightarrow u_{d}=0\quad\text{ on }\Gamma.$
Next,
[…]
and
$2[(D\bm{u})\bm{n}]_{\tau}=2(D\bm{u})\bm{n}-2((D\bm{u})\bm{n}\cdot\bm{n})\bm{n}=-\begin{pmatrix}\partial_{d}\bm{u}^{\prime}\\ 0\end{pmatrix}.$
Therefore, we can rewrite (14) as follows
(16) $(\lambda+\alpha)\bm{u}^{\prime}-\partial_{d}\bm{u}^{\prime}$ $=\bm{\phi}\quad\text{ on }\Gamma.$
Now, we aplly $\partial_{d}\text{div}^{\prime}$ to the first $(d-1)$ equations of (11) and $-\Delta^{\prime}$ to the last one; adding these equations together we obtain
$(\lambda-\Delta)$ $\left[\partial_{d}(\partial_{1}u_{1}+\partial_{2}u_{2}+\cdots+\partial_{d-1}u_{d-1})-\Delta^{\prime}u_{d}\right]$
$\qquad\quad+\left[\partial_{d}(\partial_{1}^{2}\pi+\partial_{2}^{2}\pi+\cdots+\partial_{d-1}^{2}\pi)-\Delta^{\prime}\partial_{d}\pi\right]=0.$
The pressure term clearly vanishes and in the first bracket we use $\text{div}\,\bm{u}=0\Leftrightarrow\text{div}^{\prime}\,\bm{u}^{\prime}=-\partial_{d}u_{d}$ to get
$-(\lambda-\Delta)\Delta u_{d}$ $=0\quad\text{ in }\Omega,$
$u_{d}$ $=0\quad\text{ on }\Gamma.$
Finally, we deal with the boundary system. We apply $-\text{div}^{\prime}$ to (16) to get
$-(\lambda+\alpha)\,\text{div}^{\prime}\,\bm{u}^{\prime}+\partial_{d}\text{div}^{\prime}\,\bm{u}^{\prime}=-\text{div}^{\prime}\,\bm{\phi},$
and thanks to $\text{div}^{\prime}\,\bm{u}^{\prime}=-\partial_{d}u_{d}$ we achieve
$(\lambda+\alpha-\partial_{d})\partial_{d}u_{d}=-\text{div}^{\prime}\,\bm{\phi}\quad\text{ on }\Gamma.$
Therefore, we reduced the whole problem to the question of solving the following system for $u_{d}$
(17) $\begin{split}-(\lambda-\Delta)\Delta u_{d}&=0\quad\text{ in }\Omega,\\ (\lambda+\alpha-\partial_{d})\partial_{d}u_{d}&=-\text{div}^{\prime}\,\bm{\phi}\quad\text{ on }\Gamma,\\ u_{d}&=0\quad\text{ on }\Gamma.\end{split}$
3.2. Fourier transform
We solve (17) by introducing the Fourier transform $\mathcal{F}_{x^{\prime}\to\xi}$. We need to deal with the system
(18) $\begin{split}(\lambda+|\xi|^{2}-\partial_{d}^{2})(|\xi|^{2}-\partial_{d}^{2})\widehat{u_{d}}&=0\quad\text{ for }x_{d}>0,\\ (\lambda+\alpha-\partial_{d})\partial_{d}\widehat{u_{d}}&=-i\xi\cdot\widehat{\bm{\phi}}\quad\text{ for }x_{d}=0,\\ \widehat{u_{d}}(\xi,0)&=0,\end{split}$
for any $\xi\in\mathbb{R}^{d-1}$. We see that characteristic roots are
±$\pm\sqrt{\lambda+|\xi|^{2}}\quad\text{ and }\pm|\xi|\,;$
certainly, only the terms with the minus sign are relevant. It will be useful to introduce the function
(19) $m_{0}(\lambda,\xi,x_{d})=\frac{e^{-x_{d}\sqrt{\lambda+|\xi|^{2}}}-e^{-x_{d}|\xi|}}{\lambda+(\lambda+\alpha)(\sqrt{\lambda+|\xi|^{2}}-|\xi|)}\,,$
which is nothing else than the fundamental solution to (18), i.e.
$(\lambda+|\xi|^{2}-\partial_{d}^{2})(|\xi|^{2}-\partial_{d}^{2})m_{0}$ $=0\quad\text{ for }x_{d}>0,$
$(\lambda+\alpha-\partial_{d})\partial_{d}m_{0}$ $=-1\quad\text{ for }x_{d}=0,$
$m_{0}(\lambda,\xi,0)$ $=0.$
The solution for the last component is thus
$\widehat{u_{d}}(\xi,x_{d})=i\xi\cdot m_{0}(\lambda,\xi,x_{d})\widehat{\bm{\phi}}(\xi).$
Now, we need to express $\widehat{\pi}$ and $\widehat{\bm{u}^{\prime}}$ using $m_{0}$. We start with the pressure. Let us apply $\text{div}^{\prime}$ to the first $(d-1)$ equations of (11); we get
$\Delta^{\prime}\pi=-\text{div}^{\prime}(\lambda-\Delta)\bm{u}^{\prime}\quad\text{ in }\Omega.$
Because of $\text{div}^{\prime}\,\bm{u}^{\prime}=-\partial_{d}u_{d}$ we rewrite it in the form
$\Delta^{\prime}\pi=(\lambda-\Delta)\partial_{d}u_{d}\quad\text{ in }\Omega$
and use the Fourier transform to achieve
$\widehat{\pi}=-\frac{1}{|\xi|^{2}}(\lambda+|\xi|^{2}-\partial_{d}^{2})\partial_{d}\widehat{u_{d}}.$
In other words
$\widehat{\pi}(\xi,x_{d})$ $=-(\lambda+|\xi|^{2}-\partial_{d}^{2})\partial_{d}m_{0}(\lambda,\xi,x_{d})\frac{i\xi}{|\xi|^{2}}\cdot\widehat{\bm{\phi}}(\xi)$
$=-\frac{\lambda e^{-x_{d}|\xi|}}{\lambda+(\lambda+\alpha)(\sqrt{\lambda+|\xi|^{2}}-|\xi|)}\frac{i\xi}{|\xi|}\cdot\widehat{\bm{\phi}}(\xi).$
Up to now, everything was dimension-independent. At this point, we need to express the function $\bm{u}^{\prime}$. If $d=2$, then everything is simpler. From (12) we easily find that
$\widehat{u_{1}}(\xi,x_{d})=-\partial_{2}m_{0}(\lambda,\xi,x_{2})\widehat{\bm{\phi}}(\xi).$
In the general case $d>2$ we start in the same way, i.e. we take the Fourier transform of (12) and express our unknown vector field
$i\xi\cdot\widehat{\bm{u}^{\prime}}+\partial_{d}\widehat{u_{d}}=0\Leftrightarrow\widehat{\bm{u}^{\prime}}=\frac{i\xi}{|\xi|^{2}}\partial_{d}\widehat{u_{d}},$
which yields
$\widehat{\bm{u}^{\prime}}(\xi,x_{d})=-\partial_{d}m_{0}(\lambda,\xi,x_{d})\frac{\xi\otimes\xi}{|\xi|^{2}}\widehat{\bm{\phi}}(\xi)\,.$
This yields the solution to the Stokes system, but our condition (14) is not satisfied (in general). We need to add some correcting term $\bm{v}=(\bm{v}^{\prime},0)$. To this purpose, consider the following problem
$(\lambda-\Delta)\bm{v}^{\prime}$ $=0\quad\text{ in }\Omega,$
$\text{div}\,\bm{v}^{\prime}$ $=0\quad\text{ in }\Omega,$
$(\lambda+\alpha)\bm{v}^{\prime}-\partial_{d}\bm{v}^{\prime}$ $=\bm{h}\quad\text{ on }\Gamma,$
where
$\widehat{\bm{h}}=\widehat{\bm{\phi}}-\frac{\xi\otimes\xi}{|\xi|^{2}}\widehat{\bm{\phi}}.$
Now, by the Fourier transform we get
$(\lambda+|\xi|^{2}-\partial_{d}^{2})\widehat{\bm{v}^{\prime}}$ $=0\quad\text{ in }\Omega,$
$(\lambda+\alpha-\partial_{d})\widehat{\bm{v}^{\prime}}$ $=\widehat{\bm{h}}\quad\text{ on }\Gamma.$
Characteristic roots are now just ±$\pm\sqrt{\lambda+|\xi|^{2}}$, and thus
$\widehat{\bm{v}^{\prime}}(\xi,x_{d})=\bm{b}(\xi)e^{-x_{d}\sqrt{\lambda+|\xi|^{2}}}\,.$
From the boundary condition we find
$\bm{b}(\xi)=\frac{1}{\lambda+\alpha+\sqrt{\lambda+|\xi|^{2}}}\widehat{\bm{h}}(\xi).$
Together,
$\widehat{\bm{u}^{\prime}}(\xi,x_{d})$ $=-\partial_{d}m_{0}(\lambda,\xi,x_{d})\frac{\xi\otimes\xi}{|\xi|^{2}}\widehat{\bm{\phi}}(\xi)$
$+\frac{1}{\lambda+\alpha+\sqrt{\lambda+|\xi|^{2}}}\left(\text{Id}_{d-1}\,-\frac{\xi\otimes\xi}{|\xi|^{2}}\right)e^{-x_{d}\sqrt{\lambda+|\xi|^{2}}}\widehat{\bm{\phi}}(\xi)\,.$
To conclude this section, we summarize that for $d\geq 2$, the solution to (11-14) can be expressed as
(20) $\widehat{\bm{u}^{\prime}}(\xi,x_{d})$ $=-\partial_{d}m_{0}(\lambda,\xi,x_{d})\frac{\xi\otimes\xi}{|\xi|^{2}}\widehat{\bm{\phi}}(\xi)$
$+\frac{1}{\lambda+\alpha+\sqrt{\lambda+|\xi|^{2}}}\left(\text{Id}_{d-1}\,-\frac{\xi\otimes\xi}{|\xi|^{2}}\right)e^{-x_{d}\sqrt{\lambda+|\xi|^{2}}}\widehat{\bm{\phi}}(\xi),$
$\widehat{u_{d}}(\xi,x_{d})$ $=i\xi\cdot m_{0}(\lambda,\xi,x_{d})\widehat{\bm{\phi}}(\xi),$
$\widehat{\pi}(\xi,x_{d})$ $=-\frac{\lambda e^{-x_{d}|\xi|}}{\lambda+(\lambda+\alpha)(\sqrt{\lambda+|\xi|^{2}}-|\xi|)}\frac{i\xi}{|\xi|}\cdot\widehat{\bm{\phi}}(\xi),$
where $m_{0}$ was defined in (19). In the case of $d=2$ the second part of the first right-hand side vanishes.
4. Multipliers and $L^{p}$-bounds
The aim of this section is to establish $L^{p}$-estimates of the solution obtained above. To begin with, we note that $m_{0}$ only depends on $z=|\xi|$. Setting also for simplicity $y=x_{d}$, with a slight abuse of notation we can write
(21) $m_{0}=m_{0}(\lambda,z,y)$ $=\frac{\omega(\lambda,z)+z}{\alpha+\lambda+\omega(\lambda,z)+z}\cdot\frac{e^{-y\omega(\lambda,z)}-e^{-yz}}{\lambda},$
where $\omega(\lambda,z)=\sqrt{\lambda+z^{2}}$. From relations (20) it follows that
(22) $\widehat{\bm{u}^{\prime}}(\xi,y)$ $=m_{2}(\lambda,|\xi|,y)\frac{\xi\otimes\xi}{|\xi|^{2}}\widehat{\bm{\phi}}(\xi)+m_{3}(\lambda,|\xi|,y)\left(\text{Id}_{d-1}-\frac{\xi\otimes\xi}{|\xi|^{2}}\right)\widehat{\bm{\phi}}(\xi),$
$\widehat{u_{d}}(\xi,y)$ $=m_{1}(\lambda,|\xi|,y)\frac{i\xi}{|\xi|}\cdot\widehat{\bm{\phi}}(\xi),$
$\widehat{\pi}(\xi,y)$ $=m_{4}(\lambda,|\xi|,y)\frac{i\xi}{|\xi|}\cdot\widehat{\bm{\phi}}(\xi),$
where we have set
(23) $\begin{split}m_{1}(\lambda,z,y)&=zm_{0}(\lambda,z,y),\\ m_{2}(\lambda,z,y)&=-\partial_{y}m_{0}(\lambda,z,y),\\ m_{3}(\lambda,z,y)&=\frac{1}{\alpha+\lambda+\omega(\lambda,z)}\,e^{-y\omega(\lambda,z)},\\ m_{4}(\lambda,z,y)&=-\frac{\omega(\lambda,z)+z}{\alpha+\lambda+\omega(\lambda,z)+z}\,e^{-yz}.\end{split}$
As the mutlipliers $\xi/|\xi|$ and $\xi\otimes\xi/|\xi|^{2}$ are easily seen to satisfy the Mikhlin condition (28), it is enough to treat just the terms $m_{i}$, $i=1,\dots,4$. Denote
(24) $\Sigma_{\theta}=\{z\in\mathbb{C}\setminus\{0\};\ |\arg z|<\theta\},$
Let $\mathcal{H}^{\infty}(\Sigma_{\theta})$ be the set of all all bounded, analytic functions in $\Sigma_{\theta}$. We will use repeatedly the following estimate, which follows easily from the $\mathcal{H}^{\infty}$-calculus for the operator $(-\Delta)^{1/2}$ on $L^{p}$, see e.g. [13]. For reader’s convenience, we provide a short proof based on the classical Mikhlin’s theorem.
Lemma 4.1.
Let $\widehat{\bm{u}}(\xi,y)=m(|\xi|,y)\widehat{\bm{\phi}}(\xi)$, where $m(\cdot,y)\in\mathcal{H}^{\infty}(\Sigma_{\theta})$ for each $y\geq 0$ fixed. Set
(25) $k(y)=\sup_{z\in\mathcal{H}^{\infty}(\Sigma_{\theta})}|m(z,y)|\,,\qquad y\geq 0\,.$
Then one has the estimate
(26) $\lVert\bm{u}\rVert_{L^{p}(\Gamma)}+\lVert\bm{u}\rVert_{L^{p}(\Omega)}\leq C_{q,d}\,\big{(}k(0)+\lVert k\rVert_{L^{p}(0,\infty)}\big{)}\lVert\bm{\phi}\rVert_{L^{p}(\mathbb{R}^{d-1})}.$
Furthermore, if under these assumptions $\widehat{\bm{u}}(\xi,y)=m(|\xi|,y)\widehat{\bm{v}}(\xi,y)$ for any $y>0$ fixed, then
(27) $\lVert\bm{u}\rVert_{L^{p}(\Omega)}\leq C_{q,d}\lVert k\rVert_{L^{\infty}(0,\infty)}\lVert\bm{v}\rVert_{L^{p}(\Omega)}.$
Proof.
Let $y\geq 0$ be fixed. We want to verify the Mikhlin condition
(28) $|\xi|^{k}|\nabla^{k}m(|\xi|,y)|\leq C_{k}(y),\qquad\xi\neq 0,\ k\geq 0,$
see e.g. [11, Theorem 5.2.7]. It follows once we prove
$|s^{k}m^{(k)}(s,y)|\leq C(y)\,,\qquad s\in(0,\infty).$
Let $s\in(0,\infty)$ be fixed, and let $\Gamma_{s}$ be the largest circle centered in $s$, and contained in $\Sigma_{\theta}$. By Cauchy’s theorem
$s^{k}m^{(k)}(s,y)=\frac{1}{2\pi i}\int_{\Gamma_{s}}\frac{s^{k}m(z,y)}{(z-s)^{k+1}}\,{\rm d}z.$
Now it is easy to see that the right-hand side has a bound only depending on $\theta$ and the supremum of $m(\cdot,y)$.
As the Mikhlin condition is satisfied, we see that
$\lVert\bm{u}(\cdot,y)\rVert_{L^{p}(\mathbb{R}^{d-1})}\leq C_{q,d}\,k(y)\lVert\bm{\phi}\rVert_{L^{p}(\mathbb{R}^{d-1})}$
holds. The conclusion is immediate by Fubini’s theorem. ∎
The following two lemmas are the main technical estimates of the paper.
Lemma 4.2.
Let $\theta\in(0,\pi/4)$ be fixed. Then there exist constants $C$, $c>0$ and $\delta>0$ such that
(29) $\mathfrak{Re}\,\omega(\lambda,z)$ $\geq c(z+\sqrt{|\lambda|}),$
(30) $\left|\frac{z\left(e^{-y\omega(\lambda,z)}-e^{-yz}\right)}{\omega(\lambda,z)-z}\right|$ $\leq\frac{Ce^{-\delta y|z|}}{1+y\sqrt{|\lambda|}}\,,$
(31) $\frac{1}{|\alpha+\lambda+z+\omega(\lambda,z)|}$ $\leq\frac{C}{\alpha+|\lambda|}\,,$
(32) $\frac{1}{|\alpha+\lambda+z+\omega(\lambda,z)|}$ $\leq\frac{C}{\alpha+|z|^{2}}\,,$
for all $z\in\Sigma_{\theta}$, $y\geq 0$ and $\lambda\in\Sigma_{\pi/2}$.
Proof.
For (29) see Lemma 5.2(b), (c) in [15], (30) follows from Lemma 5.3(b) therein. To show (31), we proceed as follows.
Since $\lambda\in\Sigma_{\pi/2}$ and $z\in\Sigma_{\theta}$ for some $\theta<\pi/4$, we have $\lambda+z^{2}\in\Sigma_{\pi/2}$. Hence $\omega(z)\in\Sigma_{\pi/4}$ and $z+\omega(z)\in\Sigma_{\pi/4}$. As $\frac{1}{\alpha+\lambda}\in\Sigma_{\pi/2}$, we get $\frac{z+\omega(z)}{\alpha+\lambda}\in\Sigma_{3\pi/4}$, which means that this value is strictly away from $-1$, in particular
$\Big{|}1+\frac{z+\omega(z)}{\alpha+\lambda}\Big{|}\geq C$
holds for some $C>0$. Finally,
$\Big{|}\alpha+\lambda+z+\omega(z)\Big{|}$ $\geq C|\alpha+\lambda|=C\sqrt{(\alpha+\mathfrak{Re}\,\lambda)^{2}+\mathfrak{Im}^{2}\,\lambda}$
$\geq C\sqrt{\alpha^{2}+|\lambda|^{2}}\geq C(\alpha+|\lambda|),$
where we used that $\alpha$ and $\mathfrak{Re}\,\lambda$ are positive.
Finally, as $z-z^{2}+\lambda+\omega(z)\in\Sigma_{\pi/2}$ and $\alpha+z^{2}\in\Sigma_{2\theta}$, we have
$\Big{|}1+\frac{z-z^{2}+\lambda+\omega(z)}{\alpha+z^{2}}\Big{|}\geq C,$
and the (32) follows as before. ∎
Lemma 4.3.
Let $\theta\in(0,\pi/4)$ be fixed. Then there exist constant $C>0$ and $\delta>0$ such that for all $z\in\Sigma_{\theta}$, $y\geq 0$ and $\lambda\in\Sigma_{\pi/2}$
(33) $|m_{i}(\lambda,z,y)|$ $\leq\frac{Ce^{-\delta y|z|}}{(1+y\sqrt{|\lambda|})(\alpha+|\lambda|)}\,,$
(34) $|zm_{i}(\lambda,z,y)|+|\partial_{y}m_{i}(\lambda,z,y)|$ $\leq\frac{Ce^{-\delta y|z|}}{(1+y\sqrt{|\lambda|})\sqrt{\alpha+|\lambda|}}\,,$
for $i=1$, $2$, $3$ and
(35) $|m_{4}(\lambda,z,y)|\leq Ce^{-\delta y|z|}\,.$
Proof.
We will first show the bounds for $m_{i}$ and $zm_{i}$, for all $i$. The bound for $\partial_{y}m_{i}$ will be a simple consequence of these.
Ad $m_{1}$. Using (30) and (31) we find
$|m_{1}(\lambda,z,y)|$ $=|zm_{0}(\lambda,z,y)|\leq\frac{C}{|\lambda|}\cdot\frac{e^{-\delta|z|y}}{1+y\sqrt{|\lambda|}}\Big{|}\frac{(\omega(z)-z)(\omega(z)+z)}{\alpha+\lambda+z+\omega(z)}\Big{|}$
$\leq C\frac{e^{-\delta|z|y}}{1+y\sqrt{|\lambda|}}\cdot\frac{1}{|\alpha+\lambda+z+\omega(z)|}$
$\leq\frac{C}{(1+y\sqrt{|\lambda|})(\alpha+|\lambda|)}e^{-\delta|z|y}.$
Similarly, using (31) and (32) together, we get
$|zm_{1}(\lambda,z,y)|$ $\leq C\frac{e^{-\delta|z|y}}{1+y\sqrt{|\lambda|}}\cdot\frac{|z|}{|\alpha+\lambda+z+\omega(z)|}$
$\leq C\frac{e^{-\delta|z|y}}{1+y\sqrt{|\lambda|}}\cdot\frac{1}{\sqrt{\alpha+|\lambda|}}\cdot\frac{|z|}{\sqrt{\alpha+|z|^{2}}}\,,$
from which the estimate for $zm_{1}$ follows.
Ad $m_{3}$. Due to (29) we see that
$|e^{-y\omega(z)}|\leq e^{-cy(|z|+\sqrt{|\lambda|})}=e^{-cy\sqrt{|\lambda|}}\cdot e^{-cy|z|}\leq\frac{C^{\prime}}{1+cy\sqrt{|\lambda|}}e^{-\delta y|z|}\,,$
where $(1+cy\sqrt{|\lambda|})e^{-cy\sqrt{|\lambda|}}\leq 1+\frac{1}{e}=C^{\prime}$ was used. Thanks to an obvious analogue of (31), we obtain
$|m_{3}(\lambda,z,y)|$ $=\Big{|}\frac{1}{\alpha+\lambda+\omega(z)}e^{-y\omega(z)}\Big{|}\leq\frac{C^{\prime}}{|\alpha+\lambda+\omega(z)|}\cdot\frac{e^{-\delta y|z|}}{1+cy\sqrt{|\lambda|}}$
$\leq\frac{C}{(1+cy\sqrt{|\lambda|})(\alpha+|\lambda|)}e^{-\delta y|z|}.$
Next, using a modification of both (31) and (32), we find
$|zm_{3}(\lambda,z,y)|$ $\leq C^{\prime}\frac{|z|}{|\alpha+\lambda+\omega(z)|}\cdot\frac{e^{-\delta y|z|}}{1+cy\sqrt{|\lambda|}}$
$\leq\frac{C}{(1+cy\sqrt{|\lambda|})\sqrt{\alpha+|\lambda|}}e^{-\delta y|z|}.$
Ad $m_{2}$. We have
$m_{2}(\lambda,z,y)$ $=\partial_{y}m_{0}(\lambda,z,y)$
$=\frac{1}{\lambda}\cdot\frac{z+\omega(z)}{\alpha+\lambda+z+\omega(z)}\cdot\frac{\partial}{\partial y}(e^{-y\omega(z)}-e^{-yz})$
$=\frac{1}{\lambda}\cdot\frac{z+\omega(z)}{\alpha+\lambda+z+\omega(z)}\cdot(-\omega(z)e^{-y\omega(z)}+ze^{-yz}\pm ze^{-y\omega(z)})$
$=\frac{z+\omega(z)}{\alpha+\lambda+z+\omega(z)}\cdot\frac{z(e^{-yz}-e^{-y\omega(z)})+(z-\omega(z))e^{-y\omega(z)}}{\lambda}$
$=-m_{1}(\lambda,z,y)-\frac{1}{\alpha+\lambda+z+\omega(z)}e^{-y\omega(z)}\,.$
We have already shown that the first term is bounded, the second one can be estimated in the same way as $m_{3}(\lambda,z,y)$. Similarly,
$z\partial_{y}m_{0}(\lambda,z,y)$ $=-zm_{1}(\lambda,z,y)-\frac{z}{\alpha+\lambda+z+\omega(z)}e^{-y\omega(z)}$
is bounded as both terms are bounded due to the previous parts.
It remains to show part of (32), i.e. estimate of $\partial_{y}m_{i}$ for all $i$. For $i=1$, we have
$\partial_{y}m_{1}(\lambda,z,y)=zm_{2}(\lambda,z,y),$
hence, the bound of this follows from the already proven estimate. Next, for $i=3$, we get
$\partial_{y}m_{3}(\lambda,z,y)$ $=-\frac{\omega(z)}{\alpha+\lambda+\omega(z)}e^{-y\omega(z)}.$
There obviously holds $|\omega(z)|\leq C(\sqrt{|\lambda|}+|z|)$. Hence, the boundedness of $\partial_{y}m_{3}$ follows from already established inequalities for $i=3$. And finally, for $i=2$, we can compute that
$\partial_{y}m_{2}(\lambda,z,y)$ $=-\partial_{y}m_{1}(\lambda,z,y)+\frac{\omega(z)}{\alpha+\lambda+z+\omega(z)}e^{-y\omega(z)}.$
Invoking (32) for $i=1$, $3$, the estimate of $m_{2}$ is complete.
Concerning $m_{4}$, it is enough to invoke (23) and note that $|e^{-yz}|=e^{-(cos\theta)y}$ for $z\in\Sigma_{\theta}$. ∎
We can now state and prove the following key result.
Lemma 4.4.
The solution $\bm{u}$ to (1114) satifies the estimates
(36) $\lVert\bm{u}\rVert_{L^{p}(\Gamma)}+\lVert\bm{u}\rVert_{L^{p}(\Omega)}$ $\leq\frac{K_{1}}{|\lambda|}\lVert\bm{\phi}\rVert_{L^{p}(\Gamma)},$
(37) $\lVert\nabla\bm{u}\rVert_{L^{p}(\Omega)}$ $\leq\frac{K_{2}}{\sqrt{|\lambda|}}\lVert\bm{\phi}\rVert_{L^{p}(\Gamma)}.$
Here the constants $K_{1}$ and $K_{2}$ depend on $p$ and $d$, but are independent of $\lambda$.
Proof.
In view of (22) and Lemma 4.1, we have
[…]
where
$k(\lambda,y)=\max_{i=1,2,3}\sup_{z\in\Sigma_{\theta}}|\lambda|m_{i}(\lambda,z,y).$
By Lemma 4.3 we thus obtain
$K_{1}\leq\frac{c|\lambda|}{\alpha+|\lambda|}\left[1+\left(\int_{0}^{\infty}\frac{{\rm d}y}{(1+y\sqrt{|\lambda|})^{p}}\right)^{1/p}\right]\leq c\frac{|\lambda|^{1-1/2p}}{\alpha+|\lambda|}\,.$
This is bounded independently of $\lambda$ in a given range. Estimate of $K_{2}$ is obtained similarly. ∎
In the Hilbert case $p=2$, one obtains the estimate of $\bm{u}\in W^{s,2}$ directly.
Lemma 4.5.
The solution $\bm{u}$ to (1114) satifies the estimates
(38) $c\lVert\bm{\phi}\rVert_{W^{s-1/2,2}(\Gamma)}\leq\lVert\nabla\bm{u}\rVert_{W^{s,2}(\Omega)}\leq C\lVert\bm{\phi}\rVert_{W^{s-1/2,2}(\Gamma)}$
for any $s\in[1/2,3/2)$. In particular, $\bm{u}\in W^{3/2,2}(\Omega)$ if (and only if) $\bm{\phi}\in L^{2}(\Gamma)$.
Proof.
We will sketch the proof for $d=2$, the general case being similar. The key point is that since $p=2$, we can use the Fourier transform to characterize the Sobolev norm of $\phi$ as
$\lVert\phi\rVert_{W^{s-1/2,2}(\Gamma)}^{2}=\int_{\mathbb{R}}(1+|\xi|^{2})^{s-1/2}|\widehat{\phi}(\xi)|^{2}\,{\rm d}\xi\,,$
which we want to relate to the velocity gradient norm
$\lVert\nabla\bm{u}\rVert_{W^{s,2}(\Omega)}^{2}=\int_{\mathbb{R}^{2}}(1+\xi^{2}+\eta^{2})^{s}|\mathcal{F}(\nabla\bm{u})(\xi,\eta)|^{2}\,{\rm d}\xi{\rm d}\eta\,,$
where $\mathcal{F}$ denotes the full Fourier transform. We will also frequently use the symbol $A\simeq B$ as a shorthand for $A\leq c_{1}B$ and $B\leq c_{2}A$, where $c_{1}$, $c_{2}>0$ are some irrelevant constants, and $A\lesssim B$ for a one-sided estimate.
Since $d=2$, we have $\bm{u}=(u_{1},u_{2})$, and relations (21), (22) lead to formulas for the partial Fourier transform of ^$\widehat{\nabla}=(i\xi,\partial_{y})$
$\widehat{\nabla u_{1}}(\xi,y)$ $=-\big{(}i\xi\partial_{y}m_{0}(z,y),\partial^{2}_{y}m_{0}(z,y)\big{)}\widehat{\phi}(\xi),$
$\widehat{\nabla u_{2}}(\xi,y)$ $=\big{(}i\xi zm_{0}(z,y),z\partial_{y}m_{0}(z,y)\big{)}\frac{i\xi}{|\xi|}\widehat{\phi}(\xi),$
where $z=|\xi|$ and
$m_{0}(z,y)=R(z)\big{(}e^{-\omega(z)y}-e^{-zy}\big{)}.$
Recall that $\omega(z)=\sqrt{\lambda+z^{2}}$; for simplicity, we omit the dependence on $\lambda$, which is now fixed. Note that likewise $R(z)$ is not relevant as $|R(z)|\simeq 1$.
It will be useful to set
$Z_{j}(z,\eta)=\mathcal{F}_{y\to\eta}[\partial^{j}_{y}(e^{-\omega(z)y}-e^{-zy})],\quad j=0,1,2;$
even extension with respect to $y$ is used before applying the transform. The full Fourier transform of $\nabla\bm{u}$ now consists of
$\mathcal{F}(\nabla u_{1})(\xi,\eta)$ $=-\big{(}i\xi Z_{1}(z,\eta),Z_{2}(z,\eta)\big{)}R(z)\widehat{\phi}(\xi),$
$\mathcal{F}(\nabla u_{2})(\xi,\eta)$ $=\big{(}i\xi zZ_{0}(z,\eta),zZ_{1}(z,\eta)\big{)}R(z)\widehat{\phi}(\xi).$
The key step is to evaluate and estimate $Z_{j}(z,\eta)$. In view of the formula
$\mathcal{F}_{y\to\eta}[\exp(-a|y|)]=\frac{a}{a^{2}+\eta^{2}}\,,\quad\mathfrak{Re}\,a>0,$
we obtain
$Z_{0}(z,\eta)$ $=\frac{\omega(z)}{\omega^{2}(z)+\eta^{2}}-\frac{z}{z^{2}+\eta^{2}}=\frac{\lambda}{\omega(z)+z}\cdot\frac{\eta^{2}-z\omega(z)}{(\omega^{2}(z)+\eta^{2})(z^{2}+\eta^{2})}\,,$
$Z_{1}(z,\eta)$ $=\frac{z^{2}}{z^{2}+\eta^{2}}-\frac{\omega^{2}(z)}{\omega^{2}(z)+\eta^{2}}=\frac{\lambda\eta^{2}}{(\omega^{2}(z)+\eta^{2})(z^{2}+\eta^{2})}\,,$
$Z_{2}(z,\eta)$ $=\frac{\omega^{3}(z)}{\omega^{2}(z)+\eta^{2}}-\frac{z^{3}}{z^{2}+\eta^{2}}$
$=\frac{\lambda(\eta^{2}(z^{2}+z\omega(z)+\omega^{2}(z))+z^{2}\omega^{2}(z)))}{(z+\omega(z))(\omega^{2}(z)+\eta^{2})(z^{2}+\eta^{2})}\,.$
Recalling now that $\mathfrak{Re}\,\lambda>0$ is fixed, it is straighforward to deduce that $\omega(z)\simeq 1+z$, $\omega^{2}(z)\simeq 1+z^{2}$, and hence also $|\eta^{2}-z\omega(z)|\lesssim\eta^{2}+z^{2}+z$, and $|\eta^{2}(z^{2}+z\omega(z)+\omega^{2}(z))+z^{2}\omega^{2}(z)|\simeq(\eta^{2}+z^{2})(1+z^{2})$. We thus can write
$|\mathcal{F}(\nabla u_{1})(\xi,\eta)|$ $\simeq\big{(}|zZ_{1}(z,\eta)|+|Z_{2}(z,\eta)|\big{)}|\widehat{\phi}(\xi)|\simeq\frac{1+z}{1+z^{2}+\eta^{2}}|\widehat{\phi}(\xi)|,$
$|\mathcal{F}(\nabla u_{2})(\xi,\eta)|$ $\simeq\big{(}|z^{2}Z_{0}(z,\eta)|+|zZ_{1}(z,\eta)|\big{)}|\widehat{\phi}(\xi)|\lesssim\frac{1+z}{1+z^{2}+\eta^{2}}|\widehat{\phi}(\xi)|.$
This eventually leads to a key pointwise estimate
$|\mathcal{F}(\nabla\bm{u})(\xi,\eta)|\simeq\frac{(1+|\xi|^{2})^{1/2}}{1+|\xi|^{2}+\eta^{2}}|\widehat{\phi}(\xi)|.$
It is now straighforward to deduce
$\lVert\nabla\bm{u}\rVert_{W^{s,2}(\Omega)}^{2}$ $=\int_{\mathbb{R}^{2}}(1+|\xi|^{2}+\eta^{2})^{s}|\mathcal{F}(\nabla\bm{u})(\xi,\eta)|^{2}\,{\rm d}\xi{\rm d}\eta$
$\simeq\int_{\mathbb{R}}(1+|\xi|^{2})\Big{(}\int_{\mathbb{R}}(1+|\xi|^{2}+\eta^{2})^{s-2}\,{\rm d}\eta\Big{)}|\widehat{\phi}(\xi)|^{2}\,{\rm d}\xi$
$\simeq\int_{\mathbb{R}}(1+|\xi|^{2})^{s-1/2}|\widehat{\phi}(\xi)|^{2}\,{\rm d}\xi=\lVert\phi\rVert_{W^{s-1/2}(\Gamma)}^{2}.$
Here we have also used that
$\int_{\mathbb{R}}\frac{{\rm d}\eta}{(b^{2}+\eta^{2})^{\sigma}}\simeq b^{1-2\sigma},\quad\sigma>1/2.$
5. Proof of the main results
[…]
We start with formally recovering the resolvent estimate (7). Decompose the solution as $\bm{u}=\bm{u}^{1}+\bm{u}^{2}$, where $\bm{u}^{1}$ solves the same problem, but with the Dirichlet boundary condition, and $\bm{u}^{2}$ has zero right-hand side in $\Omega$ and takes care of the boundary condition. To be specific; the function $\bm{u}^{1}$ solves the system
(39) $(\lambda-\Delta)\bm{u}^{1}+\nabla\pi^{1}$ $=\bm{f}\quad\text{ in }\Omega,$
$\text{div}\,\bm{u}^{1}$ $=0\quad\text{ in }\Omega,$
$\bm{u}^{1}$ $=0\quad\text{ on }\Gamma,$
and $\bm{u}^{2}$ solves
(40) $(\lambda-\Delta)\bm{u}^{2}+\nabla\pi^{2}$ $=0\quad\text{ in }\Omega,$
$\text{div}\,\bm{u}^{2}$ $=0\quad\text{ in }\Omega,$
$\bm{u}^{2}\cdot\bm{n}$ $=0\quad\text{ on }\Gamma,$
$(\lambda+\alpha)\bm{u}^{2}+2[(D\bm{u}^{2})\bm{n}]_{\tau}$ $=\bm{\phi}\quad\text{ on }\Gamma,$
where
(41) $\bm{\phi}=\bm{h}-2[(D\bm{u}^{1})\bm{n}]_{\tau}.$
Assume that $\lambda\in\mathbb{C}$ with $|\lambda|>\omega$, where $\omega>0$ is fixed. Due to the standard theory (see e.g. [8, Theorem 1.2]) we know that
(42) $||\lambda\bm{u}^{1}||_{L^{p}(\Omega)}+||\nabla^{2}\bm{u}^{1}||_{L^{p}(\Omega)}\leq c||\bm{f}||_{L^{p}(\Omega)}\,,$
and in particular,
$||\bm{u}^{1}||_{L^{p}(\Omega)}\leq\frac{c}{|\lambda|}||\bm{f}||_{L^{p}(\Omega)}\,.$
Because of Sobolev embedding we get
$||[(D\bm{u}^{1})\bm{n}]_{\tau}||_{L^{p}(\Gamma)}\leq c||\bm{f}||_{L^{p}(\Omega)},$
and, thus, there also holds
$||\bm{\phi}||_{L^{p}(\Gamma)}$ $=||\bm{h}-2[(D\bm{u}^{1})\bm{n}]_{\tau}||_{L^{p}(\Gamma)}\leq c||(\bm{f},\bm{h})||_{{L^{p}(\Omega)}\times{L^{p}(\Gamma)}}.$
Note that $c$ remains bounded as $|\lambda|\to\infty$. Hence, it suffices to show
$||(\bm{u}^{2},\text{tr}\,\bm{u}^{2})||_{{L^{p}(\Omega)}\times{L^{p}(\Gamma)}}\leq\frac{c}{|\lambda|}||\bm{\phi}||_{L^{p}(\Gamma)}.$
[…]
(43) $\lVert\bm{u}\rVert_{W^{1,p}(\Omega)}\leq C\big{(}\lVert\bm{f}\rVert_{L^{p}(\Omega)}+\lVert\bm{h}\rVert_{L^{p}(\Gamma)}\big{)},$
where $C$ depends on $\lambda$. Let us show uniqueness; clearly, it suffices to consider $\bm{u}$ a solution to the homogeneous problem. We set $\widetilde{\bm{f}}=|\bm{u}|^{p-2}\overline{\bm{u}}$, $\widetilde{\bm{h}}=|\bm{u}|^{p-2}\overline{\bm{u}}$ and observe that $\widetilde{\bm{f}}\in L^{p^{\prime}}(\Omega)$, $\widetilde{\bm{h}}\in L^{p^{\prime}}(\Gamma)$. By the above, there is some $\widetilde{\bm{u}}\in W^{1,p^{\prime}}_{\sigma,n}(\Omega)$, a solution corresponding to these data. Testing the weak formulation for $\bm{u}$ by $\widetilde{\bm{u}}$ and vice versa, one deduces that
$0=\int_{\Omega}\widetilde{\bm{f}}\cdot\bm{u}\,{\rm d}x+\int_{\Gamma}\widetilde{\bm{h}}\cdot\bm{u}\,{\rm d}S=\lVert\bm{u}\rVert_{L^{p}(\Omega)}^{p}+\lVert\bm{u}\rVert_{L^{p}(\Gamma)}^{p}\,.$
Hence $\bm{u}=0$.
[…]
We start with part (ii). Employing again the splitting (3941), regularity of $\bm{u}^{1}$ is proven in [10] or [8]. Note that $\bm{\phi}\in B^{1-1/p,p}_{p}(\Gamma)$ by the trace theorem [3, Theorem 7.39].
Concerning $\bm{u}^{2}$, we note that for a flat boundary, we have the so-called Robin boundary condition, whence the $W^{2,p}$-regularity follows by [17] or [12].
[…]
Introduce an auxiliary function $\bm{v}=(\bm{v}^{\prime},0)$ as the solution to
$(-\delta^{2}\Delta^{\prime}-\partial_{d}^{2})\bm{v}^{\prime}$ $=0\quad\text{ in }\Omega,$
$-\partial_{d}\bm{v}^{\prime}$ $=\bm{\phi}\quad\text{ on }\Gamma,$
where $\delta>0$ is a suitably small number. We know (see e.g. [4, Theorem 14.1]) that
(44) $\lVert\nabla^{2}\bm{v}\rVert_{L^{p}(\Omega)}\leq c\lVert\bm{\phi}\rVert_{B^{1-1/p,p}_{p}(\Gamma)}\,.$
We proceed by noting that
$\widehat{\nabla^{\prime}\bm{v}}(\xi,x_{d})=\frac{\xi}{\delta|\xi|}e^{-\delta|\xi|x_{d}}\widehat{\bm{\phi}}(\xi),\quad\textrm{or}\qquad\frac{\xi}{|\xi|}\cdot\widehat{\bm{\phi}}(\xi)=\delta e^{\delta|\xi|x_{d}}\widehat{\nabla^{\prime}\bm{v}}(\xi,x_{d}),$
and hence, in view of (22), we can write
$\widehat{\pi}(\xi,y)=f(z)e^{-\eta zy}\widehat{\nabla^{\prime}\bm{v}}(\xi,y)$
Here $\eta=1-\delta$, $z=|\xi|$, $y=x_{d}$ and
$f(z)=\frac{\omega(\lambda,z)+z}{\alpha+\lambda+\omega(\lambda,z)+z}$
is bounded and holomorphic in $\Sigma_{\pi/2}$. Hence
$\widehat{\nabla\pi}(\xi,y)$ $=(i\xi,\partial_{y})\big{[}f(z)e^{-\eta zy}\widehat{\nabla^{\prime}\bm{v}}(\xi,y)\big{]}$
$=\big{[}f(z)e^{-\eta zy}\big{]}\big{(}\widehat{(\nabla^{\prime})^{2}\bm{v}}(\xi,y),(-\eta z+\partial_{y})\widehat{\nabla^{\prime}\bm{v}}(\xi,y)\big{)}.$
The trick now is to rewrite the last term as
$(-\eta z+\partial_{y})\widehat{\nabla^{\prime}\bm{v}}(\xi,y)=-\frac{i\eta\xi}{|\xi|}\cdot\widehat{(\nabla^{\prime})^{2}\bm{v}}(\xi,y)+\widehat{\partial_{y}\nabla^{\prime}\bm{v}}(\xi,y).$
[…]
$\lVert\nabla\pi\rVert_{L^{p}(\Omega)}\leq c\lVert\nabla^{2}\bm{v}\rVert_{L^{p}(\Omega)}.$
In view of (44), the conclusion follows.
It remains to show part (i). We note first that $\bm{u}^{1}\in W^{2,r}(\Omega)$, and thus $\nabla\bm{u}^{1}\in W^{1,p}(\Omega)$. We also have $\nabla\pi^{1}\in L^{r}(\Omega)$. The key step is to bound $D\bm{u}^{1}\in B^{-1/p,p}_{p}(\Gamma)$, to which we need the following auxiliary result.
Lemma 5.1.
Let $\bm{U}\in L^{p}(\Omega)$ and $\operatorname{div}\bm{U}\in L^{r}(\Omega)$, where $1/r=1/p+1/d$. Then $(\bm{U}\cdot\bm{n})_{\tau}\in B^{-1/p,p}_{p;n}(\Gamma)$.
Proof.
Integrating by parts, we write
$\int_{\Gamma}(\bm{U}\cdot\bm{n})\bm{\varphi}\,{\rm d}S=\int_{\Omega}\bm{U}\cdot\nabla\bm{\varphi}\,{\rm d}x+\int_{\Omega}(\operatorname{div}\bm{U})\bm{\varphi}\,{\rm d}x$
and note that the right-hand side can be bounded in terms of $\bm{\varphi}\in W^{1,p^{\prime}}_{n}(\Omega)$. It remains to observe that
$\gamma(W^{1,p^{\prime}}_{n}(\Omega))=B^{1/p,p^{\prime}}_{p^{\prime};n}=(B^{-1/p,p}_{p;n}(\Gamma))^{*}\,.$
Applying the above lemma to $\bm{U}=\nabla\bm{u}^{1}$, we see that $\bm{u}^{2}$ solves (4041) where $\bm{\phi}\in B^{-1/p,p}_{p;n}(\Gamma)$.
Set $\bm{v}:=\mathcal{J}_{1}^{x^{\prime}}\bm{u}^{2}$, where $\mathcal{J}_{1}^{x^{\prime}}$ is the Bessel potential, applied in the tangential directions $x^{\prime}$. Obviously, $\bm{v}$ solves (40) with the boundary data $\bm{\Phi}=\mathcal{J}_{1}^{x^{\prime}}\bm{\phi}\in B^{1-1/p,p}$. Hence $\bm{v}\in W^{2,p}(\Omega)$, and thus $\bm{u}^{2}\in W^{1,p}(\Omega)$ as required. ∎

### v3

@@
-On $L^{p}$-semigroup to Stokes equation with dynamic boundary condition in the half-space
+On $L^{p}$-semigroup to Stokes equation with dynamic slip boundary condition in the half-space
 Abstract.
-We consider evolutionary Stokes system, coupled with the so-called dynamic boundary condition, in the simple geometry of a $d$-dimensional half-space.
+We consider evolutionary Stokes system, coupled with the so-called dynamic slip boundary condition, in the simple geometry of a $d$-dimensional half-space.
 Using the Fourier transform, we obtain an explicit formula for the resolvent.
-Maximal regularity estimates and existence of analytic semigroup in the $L^{p}$-setting are then deduced using the methods of $\mathcal{H}^{\infty}$-calculus.
+Optimal regularity estimates and existence of analytic semigroup in the $L^{p}$-setting are then deduced using the methods of $\mathcal{H}^{\infty}$-calculus.
 Key words and phrases:
-Stokes equations, dynamic boundary conditions, unbounded domains, maximal regularity, analytic semigroup
+Stokes equations, dynamic slip boundary condition, unbounded domains, optimal regularity, analytic semigroup
 2000 Mathematics Subject Classification:
@@
 This paper is motivated by the system
-(1) $\partial_{t}\bm{u}-\operatorname{div}2\nu D\bm{u}+\nabla\pi$ $=\bm{f},\qquad\operatorname{div}\bm{u}=0\quad\textrm{ in $(0,T)\times\Omega$},$
-$\beta\partial_{t}\bm{u}+\alpha\bm{u}+2\nu[{(D\bm{u})}\bm{n}]_{\tau}$ $=\beta\bm{h},\qquad\bm{u}\cdot\bm{n}=0\quad\textrm{ on $(0,T)\times\Gamma$},$
+(1) $\partial_{t}\bm{u}-\operatorname{div}2\nu{\rm D}\bm{u}+\nabla\pi$ $=\bm{f},\qquad\operatorname{div}\bm{u}=0\quad\textrm{ in $(0,T)\times\Omega$},$
+$\beta\partial_{t}\bm{u}+\alpha\bm{u}+2\nu[{({\rm D}\bm{u})}\bm{n}]_{\tau}$ $=\beta\bm{h},\qquad\bm{u}\cdot\bm{n}=0\quad\textrm{ on $(0,T)\times\Gamma$},$
 where $\Omega=\{x_{d}>0\}$ is the $d$-dimensional half-space, and $\Gamma=\{x_{d}=0\}$ the corresponding boundary.
@@
 also the weak formulation (6) below.
-However, for $p\neq 2$, the $L^{p}$-theory is expectedly more difficult and up to our knowledge, is so far missing in the literature.
-In the present paper, we intend to partially amend this, with two main results being: the existence of an analytic semigroup in $L^{p}(\Omega)\times L^{p}(\Gamma)$, and the optimal $W^{2,p}$ and $W^{1,p}$-regularity estimates for the corresponding stationary (resolvent) problem.
-Although our results are formulated for the case of half-space $\Omega=\{x_{d}>0\}$ only, extension to more general domains should be possible, using standard localization and perturbation arguments.
-Note that this would require extension of our theorems to the case of functions $\bm{u}$ with non-zero divergence, which is rather straighforward, using the estimates of [8].
-The paper is organized as follows: In Section 2, we introduce function spaces and state our main results: Theorem 2.1 and Theorem 2.2, providing resolvent estimates and optimal regularity in the $L^{p}$-setting.
-In Corollary 2.1, we show the existence of an analytic semigroup.
-We also identify the domain of its generator.
-Main technical results of the paper then follow: in Section 3, we use (partial) Fourier transform, in the spirit of [8], to express the resolvent in terms of the boundary right-hand side.
-Related key estimates in $L^{p}$ spaces are then obtained in Section 4.
-Finally, proofs of the main theorems are completed in Section 5, essentially by combining results of previous section with the well-known $L^{p}$-theory for the (homogeneous) Dirichlet problem.
-Let us conclude this introduction by discussing some related, both classical and recent results.
-For the Stokes system subject to a homogeneous Dirichlet boundary conditions, the existence of $L^{p}$-analytic semigroup was proven in [10] via the theory of pseudo-differential operators.
-For a more elementary approach based on the Fourier transform and the classical multiplier theorems, see [8].
+However, the $L^{p}$-theory is expectedly more difficult if $p\neq 2$ and up to our knowledge, the problem (1) has not been treated in the literature (although a number of closely related results are available, see also the discussion below).
+In the present paper, we intend to contribute to this, with two main results being: the existence of an analytic semigroup in $L^{p}(\Omega)\times L^{p}(\Gamma)$, and the optimal $W^{2,p}$ and $W^{1,p}$-regularity estimates for the corresponding stationary (resolvent) problem.
+Our paper is organized as follows.
+In Section 2, we introduce function spaces and state two main results: Theorem 2.1 and Theorem 2.2, providing resolvent estimates of optimal regularity both for weak and strong solutions.
+In Corollary 2.1, we show the existence of an analytic semigroup; we also identify the domain of its generator.
+Main technical results of the paper follow: in Section 3, we use (partial) Fourier transform, in the spirit of [9], to express the resolvent in terms of the boundary right-hand side.
+Key estimates in $L^{p}$ spaces are then obtained in Section 4.
+Our approach remains more elementary in that we only rely on the classical Mikhlin multiplier theorem.
+One can however note that Lemma 4.1 in fact asserts the existence of $\mathcal{H}^{\infty}$-calculus for the operator $(-\Delta)^{1/2}$.
+Finally, proofs of the main theorems are completed in Section 5, essentially by combining results of Section 4 with the well-known $L^{p}$-theory for the (homogeneous) Dirichlet problem, as given e.g.
+in [10].
+Although our results are formulated for the case of half-space $\Omega=\{x_{d}>0\}$ only, extension to more general domains is straighforward, using standard localization and perturbation arguments.
+Crucially, this requires extension of the estimates to the case of functions $\bm{u}$ with non-zero divergence.
+We treat this situation explicitly in part (iii) of Theorem 2.2.
+Let us conclude the introduction by discussing some related, both classical and more recent, results.
+Concerning the Stokes system subject to a homogeneous Dirichlet boundary condition, the existence of $L^{p}$-analytic semigroup was proven in [12] via the theory of pseudo-differential operators.
+For a more elementary approach based on the Fourier transform and classical multiplier theorems, see [9].
 Recently, a different method using the maximal operator was proposed in [18]; however, the method seems to be limited to bounded domains.
@@
 A more modern approach based on the $\mathcal{H}^{\infty}$-calculus was employed here.
-A rather comprehensive functional setup for treating resolvent problems related to dynamic boundary conditions is developed in [7], see also the references therein.
+More recently, a closely related class of problems with dynamic boundary conditions was treated in [6], and maximal regularity results were obtained.
+A rather comprehensive functional setup for treating resolvent problems related to dynamic boundary conditions is described in [8]; see also the references therein.
 This method does provide sharp regularity estimates in fractional Sobolev spaces, yet direct application to Stokes problem – due to the incompressibility condition and the pressure term – does not seem viable.
@@
 For $p\in(1,\infty)$ and $k\geq 1$ integer, we denote by $L^{p}$ and $W^{k,p}$ the usual Lebesgue and Sobolev spaces.
+Negative Sobolev space $W^{-1,p}$ is defined as the completion of $L^{q}$ with respect to the norm
+$\lVert\bm{u}\rVert_{W^{-1,p}(\Omega)}=\sup\Big{\{}\int_{\Omega}\bm{u}\cdot\bm{\varphi}\,{\rm d}x;\ \lVert\bm{\varphi}\rVert_{W^{1,p^{\prime}}(\Omega)}\leq 1\Big{\}}.$
+It it well-known (see e.g. [3, Section 3.14]) that $W^{-1,p}$ can be identified with the dual to $W^{1,p^{\prime}}$.
 We also need Besov spaces $B^{s,p}_{p}$ with a general real $s$; we allow for non-integer $s$ in the spirit of [4].
@@
 (2) $(\lambda-\Delta)\bm{u}+\nabla\pi$ $=\bm{f}\quad\text{ in }\Omega,$
-(3) $\text{div}\,\bm{u}$ $=0\quad\text{ in }\Omega,$
+(3) $\operatorname{div}\bm{u}$ $=0\quad\text{ in }\Omega,$
 (4) $\bm{u}\cdot\bm{n}$ $=0\quad\text{ on }\Gamma,$
-(5) $(\lambda+\alpha)\bm{u}+2[(D\bm{u})\bm{n}]_{\tau}$ $=\bm{h}\quad\text{ on }\Gamma,$
+(5) $(\lambda+\alpha)\bm{u}+2[({\rm D}\bm{u})\bm{n}]_{\tau}$ $=\bm{h}\quad\text{ on }\Gamma,$
 where $\lambda$ is a complex number.
-Note that $\operatorname{div}2D\bm{u}=\Delta\bm{u}$.
+Note that $\operatorname{div}2{\rm D}\bm{u}=\Delta\bm{u}$.
 Multiplying (2) by $\bm{\varphi}\in C^{\infty}_{0,\sigma}(\overline{\Omega})$ and integrating by parts, one obtains the integral formulation
-(6) $\begin{gathered}\lambda\int_{\Omega}\bm{u}\cdot\bm{\varphi}\,{\rm d}x+(\lambda+\alpha)\int_{\Gamma}\bm{u}\cdot\bm{\varphi}\,{\rm d}S+\int_{\Omega}2D\bm{u}:D\bm{\varphi}\,{\rm d}x\\ =\int_{\Omega}\bm{f}\cdot\bm{\varphi}\,{\rm d}x+\int_{\Gamma}\bm{h}\cdot\bm{\varphi}\,{\rm d}S.\end{gathered}$
-This problem is easy to handle in the $L^{2}$-setting, using Korn’s inequality, see [2] or [1]. However, our objective is to obtain the $L^{p}$ theory, as stated in the following two main theorems.
+(6) $\begin{gathered}\lambda\int_{\Omega}\bm{u}\cdot\bm{\varphi}\,{\rm d}x+(\lambda+\alpha)\int_{\Gamma}\bm{u}\cdot\bm{\varphi}\,{\rm d}S+\int_{\Omega}2{\rm D}\bm{u}:{\rm D}\bm{\varphi}\,{\rm d}x\\ =\int_{\Omega}\bm{f}\cdot\bm{\varphi}\,{\rm d}x+\int_{\Gamma}\bm{h}\cdot\bm{\varphi}\,{\rm d}S.\end{gathered}$
+This problem is easy to handle in the $L^{2}$-setting, using Korn’s inequality, see [2] or [1]. However, our objective is to obtain the $L^{p}$-theory, as stated in the following two main theorems.
 Theorem 2.1.
@@
 Theorem 2.2.
-Let $\alpha\geq 0$, $p\in(1,+\infty)$ and $\lambda$ is complex with $\mathfrak{Re}\,\lambda>0$.
+Let $\alpha\geq 0$, $p\in(1,+\infty)$ and $\lambda\in\mathbb{C}$ with $\mathfrak{Re}\,\lambda>0$.
   • (i)
-    Let $\bm{f}\in L^{r}_{\sigma,n}(\Omega)$, where $r$ is given by the relation $1/r=1/p+1/d$, and let $\bm{h}\in B^{-1/p,p}_{p;n}(\Gamma)$. Then one has the estimate
-    (8) $\lVert\bm{u}\rVert_{W^{1,p}(\Omega)}\leq C_{1}\big{(}\lVert\bm{f}\rVert_{L^{r}(\Omega)}+\lVert\bm{h}\rVert_{B^{-1/p,p}_{p}(\Gamma)}\big{)}.$
+    Let $\bm{f}\in W^{-1,p}(\Omega)$, and let $\bm{h}\in B^{-1/p,p}_{p;n}(\Gamma)$. Then one has the estimate
+    (8) $\lVert\bm{u}\rVert_{W^{1,p}(\Omega)}\leq C_{1}\big{(}\lVert\bm{f}\rVert_{W^{-1,p}(\Omega)}+\lVert\bm{h}\rVert_{B^{-1/p,p}_{p}(\Gamma)}\big{)}.$
   • (ii)
-    Let $\bm{f}\in L^{p}_{\sigma,n}(\Omega)$ and $\bm{h}\in B^{1-1/p,p}_{p;n}(\Gamma)$. Then the solution to resolvent problem (2)-(5) is strong, and one has the estimate
+    Let $\bm{f}\in L^{p}_{\sigma,n}(\Omega)$ and $\bm{h}\in B^{1-1/p,p}_{p;n}(\Gamma)$. Then the solution to resolvent problem (2)–(5) is strong, and one has the estimate
     (9) $\lVert\bm{u}\rVert_{W^{2,p}(\Omega)}+\lVert\nabla\pi\rVert_{L^{p}(\Omega)}$ $\leq C_{2}\big{(}\lVert\bm{f}\rVert_{L^{p}(\Omega)}+\lVert\bm{h}\rVert_{B^{1-1/p,p}_{p}(\Gamma)}\big{)}.$
+  • (iii)
+    If (3) is generalized to $\operatorname{div}\bm{u}=g$, the estimates remain valid, with the norm of $g$ in $L^{p}(\Omega)$ or $W^{1,p}(\Omega)$ added to the right-hand side of (8) or (9), respectively.
 The constants $C_{1}$, $C_{2}$ depend on $d$, $p$ and $|\lambda|$.
@@
 3. Fundamental solution
-Here, we deal with (2)-(5) for $\bm{f}=0$, $g=0$, i.e. we consider the system
+Here, we deal with (2)–(5) for $\bm{f}=0$, i.e. we consider the system
 (11) $(\lambda-\Delta)\bm{u}+\nabla\pi$ $=0\quad\text{ in }\Omega,$
-(12) $\text{div}\,\bm{u}$ $=0\quad\text{ in }\Omega,$
+(12) $\operatorname{div}\bm{u}$ $=0\quad\text{ in }\Omega,$
 (13) $\bm{u}\cdot\bm{n}$ $=0\quad\text{ on }\Gamma,$
-(14) $(\lambda+\alpha)\bm{u}+2[(D\bm{u})\bm{n}]_{\tau}$ $=\bm{\phi}\quad\text{ on }\Gamma.$
+(14) $(\lambda+\alpha)\bm{u}+2[({\rm D}\bm{u})\bm{n}]_{\tau}$ $=\bm{\phi}\quad\text{ on }\Gamma.$
 We will (formally) solve the problem, using Fourier analysis.
 3.1. Reducing the equations
-By $\text{div}^{\prime}$, $\Delta^{\prime}$ we denote the corresponding operator with respect to $x^{\prime}$. Let us now simplify (13)-(14). First, $\bm{n}=(0,-1)$ and thus
+By $\operatorname{div}^{\prime}$, $\Delta^{\prime}$ we denote the corresponding operator with respect to $x^{\prime}$. Let us now simplify (13)–(14). First, $\bm{n}=(0,-1)$ and thus
 (15) $\bm{u}\cdot\bm{n}=0\Leftrightarrow u_{d}=0\quad\text{ on }\Gamma.$
@@
 and
-$2[(D\bm{u})\bm{n}]_{\tau}=2(D\bm{u})\bm{n}-2((D\bm{u})\bm{n}\cdot\bm{n})\bm{n}=-\begin{pmatrix}\partial_{d}\bm{u}^{\prime}\\ 0\end{pmatrix}.$
+$2[({\rm D}\bm{u})\bm{n}]_{\tau}=2({\rm D}\bm{u})\bm{n}-2(({\rm D}\bm{u})\bm{n}\cdot\bm{n})\bm{n}=-\begin{pmatrix}\partial_{d}\bm{u}^{\prime}\\ 0\end{pmatrix}.$
 Therefore, we can rewrite (14) as follows
 (16) $(\lambda+\alpha)\bm{u}^{\prime}-\partial_{d}\bm{u}^{\prime}$ $=\bm{\phi}\quad\text{ on }\Gamma.$
-Now, we aplly $\partial_{d}\text{div}^{\prime}$ to the first $(d-1)$ equations of (11) and $-\Delta^{\prime}$ to the last one; adding these equations together we obtain
+Now, we aplly $\partial_{d}\operatorname{div}^{\prime}$ to the first $(d-1)$ equations of (11) and $-\Delta^{\prime}$ to the last one; adding these equations together we obtain
 $(\lambda-\Delta)$ $\left[\partial_{d}(\partial_{1}u_{1}+\partial_{2}u_{2}+\cdots+\partial_{d-1}u_{d-1})-\Delta^{\prime}u_{d}\right]$
 $\qquad\quad+\left[\partial_{d}(\partial_{1}^{2}\pi+\partial_{2}^{2}\pi+\cdots+\partial_{d-1}^{2}\pi)-\Delta^{\prime}\partial_{d}\pi\right]=0.$
-The pressure term clearly vanishes and in the first bracket we use $\text{div}\,\bm{u}=0\Leftrightarrow\text{div}^{\prime}\,\bm{u}^{\prime}=-\partial_{d}u_{d}$ to get
+The pressure term clearly vanishes and in the first bracket we use $\operatorname{div}\bm{u}=0\Leftrightarrow\\ diver^{\prime}\bm{u}^{\prime}=-\partial_{d}u_{d}$ to get
 $-(\lambda-\Delta)\Delta u_{d}$ $=0\quad\text{ in }\Omega,$
 $u_{d}$ $=0\quad\text{ on }\Gamma.$
-Finally, we deal with the boundary system. We apply $-\text{div}^{\prime}$ to (16) to get
-$-(\lambda+\alpha)\,\text{div}^{\prime}\,\bm{u}^{\prime}+\partial_{d}\text{div}^{\prime}\,\bm{u}^{\prime}=-\text{div}^{\prime}\,\bm{\phi},$
-and thanks to $\text{div}^{\prime}\,\bm{u}^{\prime}=-\partial_{d}u_{d}$ we achieve
-$(\lambda+\alpha-\partial_{d})\partial_{d}u_{d}=-\text{div}^{\prime}\,\bm{\phi}\quad\text{ on }\Gamma.$
+Finally, we deal with the boundary system. We apply $-\operatorname{div}^{\prime}$ to (16) to get
+$-(\lambda+\alpha)\,\operatorname{div}^{\prime}\bm{u}^{\prime}+\partial_{d}\operatorname{div}^{\prime}\bm{u}^{\prime}=-\operatorname{div}^{\prime}\bm{\phi},$
+and thanks to $\operatorname{div}^{\prime}\bm{u}^{\prime}=-\partial_{d}u_{d}$ we achieve
+$(\lambda+\alpha-\partial_{d})\partial_{d}u_{d}=-\operatorname{div}^{\prime}\bm{\phi}\quad\text{ on }\Gamma.$
 Therefore, we reduced the whole problem to the question of solving the following system for $u_{d}$
-(17) $\begin{split}-(\lambda-\Delta)\Delta u_{d}&=0\quad\text{ in }\Omega,\\ (\lambda+\alpha-\partial_{d})\partial_{d}u_{d}&=-\text{div}^{\prime}\,\bm{\phi}\quad\text{ on }\Gamma,\\ u_{d}&=0\quad\text{ on }\Gamma.\end{split}$
+(17) $\begin{split}-(\lambda-\Delta)\Delta u_{d}&=0\quad\text{ in }\Omega,\\ (\lambda+\alpha-\partial_{d})\partial_{d}u_{d}&=-\operatorname{div}^{\prime}\bm{\phi}\quad\text{ on }\Gamma,\\ u_{d}&=0\quad\text{ on }\Gamma.\end{split}$
 3.2. Fourier transform
@@
 $\widehat{u_{d}}(\xi,x_{d})=i\xi\cdot m_{0}(\lambda,\xi,x_{d})\widehat{\bm{\phi}}(\xi).$
-Now, we need to express $\widehat{\pi}$ and $\widehat{\bm{u}^{\prime}}$ using $m_{0}$. We start with the pressure. Let us apply $\text{div}^{\prime}$ to the first $(d-1)$ equations of (11); we get
-$\Delta^{\prime}\pi=-\text{div}^{\prime}(\lambda-\Delta)\bm{u}^{\prime}\quad\text{ in }\Omega.$
-Because of $\text{div}^{\prime}\,\bm{u}^{\prime}=-\partial_{d}u_{d}$ we rewrite it in the form
+Now, we need to express $\widehat{\pi}$ and $\widehat{\bm{u}^{\prime}}$ using $m_{0}$.
+We start with the pressure.
+Let us apply $\operatorname{div}^{\prime}$ to the first $(d-1)$ equations of (11); we get
+$\Delta^{\prime}\pi=-\operatorname{div}^{\prime}(\lambda-\Delta)\bm{u}^{\prime}\quad\text{ in }\Omega.$
+Because of $\operatorname{div}^{\prime}\bm{u}^{\prime}=-\partial_{d}u_{d}$ we rewrite it in the form
 $\Delta^{\prime}\pi=(\lambda-\Delta)\partial_{d}u_{d}\quad\text{ in }\Omega$
@@
 $(\lambda-\Delta)\bm{v}^{\prime}$ $=0\quad\text{ in }\Omega,$
-$\text{div}\,\bm{v}^{\prime}$ $=0\quad\text{ in }\Omega,$
+$\operatorname{div}\bm{v}^{\prime}$ $=0\quad\text{ in }\Omega,$
 $(\lambda+\alpha)\bm{v}^{\prime}-\partial_{d}\bm{v}^{\prime}$ $=\bm{h}\quad\text{ on }\Gamma,$
@@
 (28) $|\xi|^{k}|\nabla^{k}m(|\xi|,y)|\leq C_{k}(y),\qquad\xi\neq 0,\ k\geq 0,$
-see e.g. [11, Theorem 5.2.7]. It follows once we prove
+see e.g. [13, Theorem 5.2.7]. It follows once we prove
 $|s^{k}m^{(k)}(s,y)|\leq C(y)\,,\qquad s\in(0,\infty).$
@@
 (39) $(\lambda-\Delta)\bm{u}^{1}+\nabla\pi^{1}$ $=\bm{f}\quad\text{ in }\Omega,$
-$\text{div}\,\bm{u}^{1}$ $=0\quad\text{ in }\Omega,$
+$\operatorname{div}\bm{u}^{1}$ $=0\quad\text{ in }\Omega,$
 $\bm{u}^{1}$ $=0\quad\text{ on }\Gamma,$
@@
 (40) $(\lambda-\Delta)\bm{u}^{2}+\nabla\pi^{2}$ $=0\quad\text{ in }\Omega,$
-$\text{div}\,\bm{u}^{2}$ $=0\quad\text{ in }\Omega,$
+$\operatorname{div}\bm{u}^{2}$ $=0\quad\text{ in }\Omega,$
 $\bm{u}^{2}\cdot\bm{n}$ $=0\quad\text{ on }\Gamma,$
-$(\lambda+\alpha)\bm{u}^{2}+2[(D\bm{u}^{2})\bm{n}]_{\tau}$ $=\bm{\phi}\quad\text{ on }\Gamma,$
+$(\lambda+\alpha)\bm{u}^{2}+2[({\rm D}\bm{u}^{2})\bm{n}]_{\tau}$ $=\bm{\phi}\quad\text{ on }\Gamma,$
 where
-(41) $\bm{\phi}=\bm{h}-2[(D\bm{u}^{1})\bm{n}]_{\tau}.$
-Assume that $\lambda\in\mathbb{C}$ with $|\lambda|>\omega$, where $\omega>0$ is fixed. Due to the standard theory (see e.g. [8, Theorem 1.2]) we know that
+(41) $\bm{\phi}=\bm{h}-2[({\rm D}\bm{u}^{1})\bm{n}]_{\tau}.$
+Assume that $\lambda\in\mathbb{C}$ with $|\lambda|>\omega$, where $\omega>0$ is fixed. Due to the standard theory (see e.g. [9, Theorem 1.2]) we know that
 (42) $||\lambda\bm{u}^{1}||_{L^{p}(\Omega)}+||\nabla^{2}\bm{u}^{1}||_{L^{p}(\Omega)}\leq c||\bm{f}||_{L^{p}(\Omega)}\,,$
@@
 Because of Sobolev embedding we get
-$||[(D\bm{u}^{1})\bm{n}]_{\tau}||_{L^{p}(\Gamma)}\leq c||\bm{f}||_{L^{p}(\Omega)},$
+$||[({\rm D}\bm{u}^{1})\bm{n}]_{\tau}||_{L^{p}(\Gamma)}\leq c||\bm{f}||_{L^{p}(\Omega)},$
 and, thus, there also holds
-$||\bm{\phi}||_{L^{p}(\Gamma)}$ $=||\bm{h}-2[(D\bm{u}^{1})\bm{n}]_{\tau}||_{L^{p}(\Gamma)}\leq c||(\bm{f},\bm{h})||_{{L^{p}(\Omega)}\times{L^{p}(\Gamma)}}.$
+$||\bm{\phi}||_{L^{p}(\Gamma)}$ $=||\bm{h}-2[({\rm D}\bm{u}^{1})\bm{n}]_{\tau}||_{L^{p}(\Gamma)}\leq c||(\bm{f},\bm{h})||_{{L^{p}(\Omega)}\times{L^{p}(\Gamma)}}.$
 Note that $c$ remains bounded as $|\lambda|\to\infty$. Hence, it suffices to show
@@
 […]
-We start with part (ii).
-Employing again the splitting (3941), regularity of $\bm{u}^{1}$ is proven in [10] or [8].
+We split the solution $\bm{u}=\bm{u}^{1}+\bm{u}^{2}$ as above, taking moreover $\operatorname{div}\bm{u}^{1}=g$ in (39). Hence, part (iii) of the Theorem is taken into account.
+Let us now consider the strong estimates, i.e.
+part (ii).
+Regularity of $\bm{u}^{2}$ follows as in [9], see also [10, Theorem IV.3.2].
 Note that $\bm{\phi}\in B^{1-1/p,p}_{p}(\Gamma)$ by the trace theorem [3, Theorem 7.39].
@@
 and hence, in view of (22), we can write
-$\widehat{\pi}(\xi,y)=f(z)e^{-\eta zy}\widehat{\nabla^{\prime}\bm{v}}(\xi,y)$
+$\widehat{\pi}(\xi,y)=f(z)e^{-\eta zy}\widehat{\nabla^{\prime}\bm{v}}(\xi,y).$
 Here $\eta=1-\delta$, $z=|\xi|$, $y=x_{d}$ and
@@
 It remains to show part (i).
-We note first that $\bm{u}^{1}\in W^{2,r}(\Omega)$, and thus $\nabla\bm{u}^{1}\in W^{1,p}(\Omega)$.
-We also have $\nabla\pi^{1}\in L^{r}(\Omega)$.
-The key step is to bound $D\bm{u}^{1}\in B^{-1/p,p}_{p}(\Gamma)$, to which we need the following auxiliary result.
+Existence and regularity of $\bm{u}^{1}$ follows from [10, Theorem IV.3.3].
+The key step is to bound ${\rm D}\bm{u}^{1}\in B^{-1/p,p}_{p}(\Gamma)$, to which we need the following auxiliary result.
 Lemma 5.1.
-Let $\bm{U}\in L^{p}(\Omega)$ and $\operatorname{div}\bm{U}\in L^{r}(\Omega)$, where $1/r=1/p+1/d$. Then $(\bm{U}\cdot\bm{n})_{\tau}\in B^{-1/p,p}_{p;n}(\Gamma)$.
+Let $\bm{U}\in L^{p}(\Omega)$ and $\operatorname{div}\bm{U}\in W^{-1,p}(\Omega)$, Then $(\bm{U}\cdot\bm{n})_{\tau}\in B^{-1/p,p}_{p;n}(\Gamma)$.
 Proof.
@@
 $\int_{\Gamma}(\bm{U}\cdot\bm{n})\bm{\varphi}\,{\rm d}S=\int_{\Omega}\bm{U}\cdot\nabla\bm{\varphi}\,{\rm d}x+\int_{\Omega}(\operatorname{div}\bm{U})\bm{\varphi}\,{\rm d}x$
-and note that the right-hand side can be bounded in terms of $\bm{\varphi}\in W^{1,p^{\prime}}_{n}(\Omega)$. It remains to observe that
-$\gamma(W^{1,p^{\prime}}_{n}(\Omega))=B^{1/p,p^{\prime}}_{p^{\prime};n}=(B^{-1/p,p}_{p;n}(\Gamma))^{*}\,.$
+and note that the right-hand side can be bounded in terms of $\bm{\varphi}\in W^{1,p^{\prime}}_{n}(\Omega)$. Denoting formally by $\gamma$ the trace operator, it remains to observe that
+$\gamma(W^{1,p^{\prime}}_{n}(\Omega))=B^{1/p,p^{\prime}}_{p^{\prime};n}(\Gamma)=(B^{-1/p,p}_{p;n}(\Gamma))^{*}\,.$
 Applying the above lemma to $\bm{U}=\nabla\bm{u}^{1}$, we see that $\bm{u}^{2}$ solves (4041) where $\bm{\phi}\in B^{-1/p,p}_{p;n}(\Gamma)$.